OpenStudy (anonymous):

f(x) =x+5/x-10 what is the inverse of f(x) and how do I determine the range of the inverse

OpenStudy (anonymous):

put y=x+5/(x-10) y=(x^2-10x+5)/(x-10) x^2-10x+5=xy-10y x^2-(10+y)x+(5+10y)=0 from here find the roots of x using quadratic formula and f^-1(y)=x so the function of y u get is the inverse function

OpenStudy (dumbcow):

Range of inverse is same as domain of original function domain is all real numbers, except x not equal to 10